# Generic rigidity and accidental modes in metal-organic frameworks

### v1

@@ Abstract
Metal–organic frameworks (MOFs) combine high porosity with structural fragility, raising important questions about their mechanical stability. We develop a rigidity-based framework in which spring networks parameterized by UFF4MOF are used to construct rigidity and dynamical matrices. Large-scale analysis of 5,682 MOFs from the CoRE 2019 database shows that most frameworks are formally over-constrained yet cluster sharply near the isostatic threshold, revealing accidental geometric modes and placing many MOFs near mechanical instability. In the representative case of UiO-66, we show that auxiliary long-range constraints introduced by tuning the neighbor cutoff lift these modes into soft, flat, finite-frequency bands. The results show that rigidity-matrix analysis can rapidly identify MOFs likely to remain mechanically stable. This near-criticality mirrors behavior known from topological mechanics and points to a deeper design principle in porous crystals.
@@ I Introduction
Metal–organic frameworks (MOFs) combine crystalline order with exceptional porosity, creating vast internal surface areas and highly tunable architectures that are promising for applications ranging from gas storage to catalysis (Fig. 1[Furukawa2010, Zhou-mofs]. These same attributes often place MOFs near the edge of mechanical [MOF-prop] stability: frameworks collapse during activation, deform under modest pressure, or exhibit soft, collective motions that couple strongly to guests molecules [ZEGGAI2025100864]. Predicting which structures are robust and which are mechanically fragile remains a central challenge [MOF-prop2, Moosavi2018ChemicalCaryatids] for design and screening. Early efforts to describe MOF mechanics include the hinge–truss model of Sarkisov [Sarkisov2014], which showed that framework flexibility can often be anticipated from network topology. Marmier and Evans [MOFflex] subsequently carried out a more mathematical, group-theoretical analysis of MOF frameworks, demonstrating that simple scalar Maxwell–Calladine counts are unreliable for MOFs. Since self stress can obscure underlying mechanisms a more robust method is need thus motivating a matrix-based approached.
First principles and classical approaches each address part of this problem. Density functional theory (DFT) captures bonding and elastic responses with high accuracy but is limited to small cells [Maurer2019]. Classical force fields and machine learned interatomic potentials (MLIPs) [Moghadam2019, Cleeton2025DeepDreamMOF, MOF-ML] extend accessible scales and enable dynamics for larger systems, but remain too computationally expensive for high throughput screening and often yield descriptive metrics without revealing the structural degrees of freedom that govern rigidity.
We take a complementary route based on rigidity theory, also known as topological constraint theory (TCT) [Thorpe1983, Jacobs1995]. In this framework, a structure’s mechanical response is governed by the balance between internal degrees of freedom and the independent constraints imposed by its bonding network. The key objects are zero modes which are motions that cost no energy and states of self stress, which are patterns of internal tension and compression that satisfy force balance. The Maxwell–Calladine [Maxwell1864, Calladine1978] relation,
@@
In this work, we operationalize this program for MOFs by building constraint-based Hamiltonians with linear and angular springs, assembling rigidity matrices for periodic crystals, computing their Maxwell–Calladine indices, and phonon dispersions. The resulting modes are then characterized using the Inverse Participation Ratio (IPR) to distinguish between localized and collective motions, which is crucial for identifying the origins of framework flexibility. Spring constants are assigned from UFF4MOF [Rappe1992, Addicoat2014, UFF-Ext] parameters, and all analyses are performed in reciprocal space to resolve wave vector-dependent zero modes. Applied to 5,682 frameworks in the CoRE 2019 [Chung2024CoREMOF] dataset, this approach reveals that most MOFs occupy a mechanically marginal regime while only a smaller subset behaves as genuinely rigid frameworks. In this way, the rigidity-matrix analysis complements first-principles benchmarks: DFT and MLIPs provide accuracy for selected targets, while the present approach provides scalable guidance and interpretable relations across thousands of candidates.
Figure 1: A: MOFs are typically obtained via solvothermal methods [solvothermal], where metal ions or clusters and multidentate organic linkers self-assemble into extended crystalline networks. B: The resulting frameworks exhibit permanent porosity, high internal surface areas, and tunable structures and functionalities. C: Many MOFs are mechanically fragile and can degrade or collapse under external stress, highlighting the importance of identifying mechanically robust frameworks. D: MOFs have been investigated for gas storage, separations, heterogeneous catalysis, and chemical sensing.
Figure 2: Zero-mode count $N_{0}$ as a function of the constraint cutoff parameter $\tau$, which controls which neighbors are included in the rigidity model. In our convention, larger $\tau$ reduces the coordination weight used by CrystalNN[crystalNN], removing constraints and increasing $N_{0}$. Each colored curve corresponds to a different MOF, and purple stars ($\ast$) mark the plateau $\tau$ values selected automatically by CrystalNN. For clarity, the UiO-66 curve is scaled by a factor of 10. The inset shows the FIGXAU framework: Color indicates the baseline bonding network, while the gray version illustrates the auxiliary bonds identified when $\tau$ is reduced.
@@ II.1 Constraint-Based Model and Parameterization
where $\mu$ indexes all independent constraints in the network, including both bond-stretching terms $(i,j)\in\mathcal{L}$ and bond-bending terms $(i,j,k)\in\mathcal{A}$. The associated extensions are $e_{ij}=\hat{\mathbf{e}}_{ij}\cdot(\mathbf{u}_{i}-\mathbf{u}_{j})$ and $e_{ijk}=\theta_{ijk}-\theta^{0}_{ijk}$. Stacking all extensions into a vector $\mathbf{e}$ and displacements into $\mathbf{u}$ defines the rigidity matrix $R$ through $\mathbf{e}=R\mathbf{u}$. In our implementation, the same $R$ used for Maxwell–Calladine counting also enters the energetic dynamical matrix once local stiffness and prestress are specified, consistent with the generalized rigidity analyis of Rocks and Mehta [rocks2024]. The potential energy and dynamical matrix then follow as:
$V=\tfrac{1}{2}\,\mathbf{u}^{T}(R^{T}KR)\,\mathbf{u},\quad D=M^{-1/2}\,R^{T}K\,R\,M^{-1/2}$ (3)
Here $M$ is the diagonal mass matrix constructed from the atomic masses $m_{i}$, and $K_{\mu}$ are the bond and angle spring constants obtained from UFF4MOF parameters, with atom types assigned according to local coordination environments. Explicitly,
$K_{ij}$ $=664.12\,\frac{Z_{i}Z_{j}}{r_{ij}^{3}},$ (4)
$K_{ijk}$ $=\frac{664.12}{r_{ij}r_{jk}}\;\frac{Z_{i}Z_{k}}{r_{ik}^{5}}\,\Big[\,3\,r_{ij}r_{jk}(1-\cos^{2}\theta_{0})-r_{ik}^{2}\cos\theta_{0}\Big],$ (5)
where $Z_{\ell}$ are the UFF effective charges. We focus on these terms as they provide the dominant constraints governing framework rigidity. The UFF equilibrium distances are used as rest lengths; across all linear bonds they exceed the geometric bond lengths by 5–7% on average (see Appendix B).
The bonding network itself is determined using the CrystalNN[crystalNN] algorithm implemented in pymatgen[pymatgen]. A tunable cutoff parameter $\tau$ controls which coordination shells are included, as shown in Fig. 2. In the main dataset analysis we adopt the plateau $\tau$ identified by CrystalNN, which provides a chemically sensible minimal network, but in Fig. 4 we explore the effect of reducing $\tau$ for UiO-66, showing that modest additions of longer-range constraints systematically lift many of the baseline zero modes into soft, finite-frequency vibrations.
@@ II.2 Phonon Dispersions and Localization Analysis
which yields the $\mathbf{k}$-dependent dynamical matrix $D(\mathbf{k})$. Diagonalizing $D(\mathbf{k})$ gives the phonon spectrum $\omega^{2}(\mathbf{k})$, which we compute along the Setyawan–Curtarolo [Setyawan_2010] high-symmetry paths in the Brillouin zone.
@@ II.3 Dataset and Implementation Details
We apply this framework to 5,682 structures from the CoRE MOF 2019 database [Chung2024CoREMOF]. Primitive cells are used whenever available to reduce redundancy. CIF [CIF_note] parsing, neighbor detection, constraint assembly, and phonon analysis were automated to enable large-scale screening. To validate our method, we benchmarked the rigidity code against crystalline silicon, shown in Appendix C, and against UiO–66, whose detailed analysis is presented in the main text. Full implementation details and all input files are provided in Appendix B and in the publicly available repository referenced therein.
@@
Figure 4: (A) Benchmark phonon dispersion calculated via GULP, reproduced from Ref. [UiO-66]. (B) Corresponding result from our rigidity-based model ($N_{0}=238$) using the default $\tau$ setting. (C) Decreasing the cutoff to $\tau=0.5$ introduces auxiliary constraints, reducing the zero mode count to $N_{0}=215$. (D) Further tuning to $\tau=0.035$ eliminates all zero modes ($N_{0}=0$). The progression from (B) to (D) demonstrates how assigning finite stiffness to long-range auxiliary bonds lifts zero modes into soft, collective bands, progressively filling the low-frequency spectrum relative to the benchmark in (A).
@@
Figure 5: Rigidity statistics across the CoRE MOF dataset. (A) Scatter plot of $\nu$ for 500 representative MOFs, with the red dashed line marking the isostatic threshold ($\nu=3$). (B) Histogram of the Maxwell–Calladine index density ($\nu/N$) for 5,182 MOFs at $\Gamma$. The majority of frameworks are formally over-constrained yet cluster sharply near the isostatic point. This concentration is dominated by accidental modes, which place many MOFs on the verge of mechanical instability.
@@ III Results
Our large scale analysis of 5,682 CoRE MOF structures shows that most frameworks are formally over-constrained. As illustrated in Fig. 5, the distribution of the Maxwell–Calladine index $\nu$ is skewed toward negative values. However, rather than exhibiting a broad continuum of rigidity, these materials cluster sharply near the isostatic threshold $\nu=3$.
This sharp concentration indicates a universal mechanical trend: the self-assembly of porous frameworks drives them toward the brink of mechanical instability. In our dataset, this marginal behavior dominated by accidental geometric modes, which are zero modes that do not appear in Maxwell–Calladine counting because the assumption of generic geometry is violated. The local geometry of the metal clusters and linkers creates alignments that render some constraints redundant, which leads to accidental zero modes despite the large number of constraints. A smaller subset of truly generic frameworks, clustered around $\nu/N_{s}\approx-1.5$, display conventional stability. This confirms that accidental modes are a pervasive feature of MOF architectures rather than isolated anomalies.
To systematically characterize these behaviors, we classify frameworks into three regimes based on the relationship between $\nu$ and the vibrational spectrum:
@@
  • Geometrically Unstable ($\nu<3$, $N_{0}>3$): Symmetry creates constraint redundancy, producing accidental nontrivial mechanisms despite being formally over constrained.
@@
Structures with $\nu>3$ are classified as under-constrained and are stabilized only be weak long-range couplings. In practical terms, generically rigid frameworks are robust to small distortions, whereas geometrically unstable and singular isostatic frameworks can be pushed between rigid and soft behavior by modest geometric changes. To illustrate the physical implications of the three regimes, we examine the representative MOFs shown in Fig. 3.
ABIXOZ represents Generic Rigidity. It is strongly over-constrained with $\nu/N=-1.89$ and contains no zero modes. Its phonon dispersion in Fig. 3(3) is smooth and well ordered, and its IPR profile shows predominantly delocalized motion. ABIXOZ therefore serves as a reference for genuine topological rigidity in which the constraint count reliably predicts mechanical behavior.
IKEBUV01 exemplifies the Singular Isostatic regime. Residing at the mechanical critical point ($\nu=3$), its stability is not generic but enforced by symmetry: it contains twelve zero modes exactly canceled by twelve states of self stress. Its phonon dispersion in Fig. 3(3) reveals low frequency asymmetries and nearly flat branches, signatures of marginal rigidity. The IPR map shows that many zero modes remain delocalized, but this balance is fragile; small geometric distortions could stabilize the system but the removal of periodic boundary conditions would likely drive the structure into a generically unstable phase.
UiO–66 lies within the Geometrically Unstable regime. Despite a Maxwell–Calladine index indicating high rigidity ($\nu/N=-1.14$), its spectrum contains 238 zero modes at $\Gamma$. These are accidental geometric modes arising from degeneracies within the MOF. The phonon dispersion in Fig. 3(3) shows bands that flatten near $\Gamma$, indicating internal degrees of freedom that permit compliant motion within an otherwise over-constrained network.
@@
To confirm the geometric origin of these modes, we systematically introduce auxiliary long-range bonds by tuning the cutoff parameter $\tau$ (Fig. 4). As $\tau$ decreases, the constraint density increases. By assigning finite stiffness to these auxiliary bonds while retaining standard UFF parameters for the primary network, we observe that the zero modes are lifted to finite frequencies, manifesting as soft, collective modes.
@@
Taken together, these case studies show that formal constraint counting alone cannot predict rigidity in porous crystalline materials. While ABIXOZ behaves as a generic over-constrained network, UiO–66 and IKEBUV01 derive their mechanical character from accidental geometric modes and constraint dependencies. Distinguishing between generic and accidental modes is therefore essential for understanding and predicting the mechanical stability of MOFs.

### v2

# Singularly isostatic and geometrically unstable rigidity of metal-organic frameworks

@@ Abstract
Metal–organic frameworks (MOFs) combine high porosity with structural fragility, raising important questions about their mechanical stability. We develop a rigidity-based formalism in which spring networks parameterized by UFF4MOF are used to construct rigidity and dynamical matrices. Large-scale analysis of 5,682 MOFs from the CoRE MOF 2019 database shows that most frameworks are formally over-constrained yet cluster sharply near the isostatic threshold, revealing accidental geometric modes and placing many MOFs near mechanical instability. In the representative case of UiO-66, we show that auxiliary long-range constraints introduced by tuning the neighbor cutoff lift these modes into soft, flat, finite-frequency bands. The results show that rigidity-matrix analysis can rapidly identify MOFs likely to remain mechanically stable. This near-criticality mirrors behavior known from topological mechanics and points to a deeper design principle in porous crystals.
@@ I Introduction
Metal–organic frameworks (MOFs) combine crystalline order with exceptional porosity, creating vast internal surface areas and highly tunable architectures that are promising for applications ranging from gas storage to catalysis (Fig. 1[7, 10]. These same attributes often place MOFs near the edge of mechanical [23] stability: frameworks collapse during activation, deform under modest pressure, or exhibit soft, collective motions that couple strongly to guest molecules [32]. Predicting which structures are robust and which are mechanically fragile remains a central challenge [2, 18] for design and screening. Early efforts to describe MOF mechanics include the hinge–truss model of Sarkisov [25], which showed that framework flexibility can often be anticipated from network topology. Marmier and Evans [14] subsequently carried out a more mathematical, group-theoretical analysis of MOF frameworks, demonstrating that simple scalar Maxwell–Calladine counts are unreliable for MOFs. Since self stress can obscure underlying mechanisms a more robust method is needed thus motivating a matrix-based approach.
First principles and classical approaches each address part of this problem. Density functional theory (DFT) captures bonding and elastic responses with high accuracy but is limited to small cells [15]. Classical force fields and machine-learned interatomic potentials (MLIPs) [17, 5, 21] extend accessible scales and enable dynamics for larger systems, but remain too computationally expensive for high-throughput screening and often yield descriptive metrics without revealing the structural degrees of freedom that govern rigidity.
An alternative to first-principle or machine-learning approaches is to construct a simplified mechanical model that captures the the essence of the material. One established approach in this direction is the Rigid Unit Model (RUM), which identifies low energy vibrational modes by treating polyhedra in crystalline solids as perfectly stiff bodies connected by flexible hinges  [8]. By searching for geometric distortions that leave these rigid units undeformed, the RUM model successfully predicts low-energy vibrational modes and flexibility windows based solely on connectivity. For disordered solids, the analytical theory of nonaffine deformation  [31] is insightful. This formalism explicitly links the vibrational spectrum to mechanical stability by showing how nonaffine displacements reduce the shear modulus, thus driving the material toward a loss of rigidity as the coordination number decreases. The affine stiffness is calculated by averaging the geometric orientations of the atomic bonds. Both of these approaches utilize geometry to gain insight.
We take a complementary route based on rigidity theory, also known as topological constraint theory (TCT) [27, 12]. In this formalism, a structure’s mechanical response is governed by the balance between internal degrees of freedom and the independent constraints imposed by its bonding network. The key objects are zero modes which are motions that cost no energy and states of self stress, which are patterns of internal tension and compression that satisfy force balance. The Maxwell–Calladine [16, 3] relation,
@@
In this work, we operationalize this program for MOFs by building constraint-based Hamiltonians with linear and angular springs, assembling rigidity matrices for periodic crystals, computing their Maxwell–Calladine indices, and phonon dispersions. The resulting modes are then characterized using the Inverse Participation Ratio (IPR) to distinguish between localized and collective motions, which is crucial for identifying the origins of framework flexibility. Spring constants are assigned from UFF4MOF [22, 1, 6] parameters, and all analyses are performed in reciprocal space to resolve wave vector-dependent zero modes. Applied to 5,682 frameworks in the CoRE 2019 [4] dataset, this approach reveals that most MOFs occupy a mechanically marginal regime while only a smaller subset behaves as genuinely rigid frameworks. In this way, the rigidity-matrix analysis complements first-principles benchmarks: DFT and MLIPs provide accuracy for selected targets, while the present approach provides scalable guidance and interpretable relations across thousands of candidates.
Figure 1: A: MOFs are typically obtained via solvothermal methods [30], where metal ions or clusters and multidentate organic linkers self-assemble into extended crystalline networks. B: The resulting frameworks exhibit permanent porosity, high internal surface areas, and tunable structures and functionalities. C: Many MOFs are mechanically fragile and can degrade or collapse under external stress, highlighting the importance of identifying mechanically robust frameworks. D: MOFs have been investigated for gas storage, separations, heterogeneous catalysis, and chemical sensing.
Figure 2: Zero-mode count $N_{0}$ as a function of the constraint cutoff parameter $\tau$, which controls which neighbors are included in the rigidity model. In our convention, larger $\tau$ reduces the coordination weight used by CrystalNN[20], removing constraints and increasing $N_{0}$. Each colored curve corresponds to a different MOF, and purple stars ($\ast$) mark the plateau $\tau$ values selected automatically by CrystalNN. For clarity, the UiO-66 curve is scaled by a factor of 10. The inset shows the FIGXAU framework: Color indicates the baseline bonding network, while the gray version illustrates the auxiliary bonds identified when $\tau$ is reduced.
@@ II.1 Constraint-Based Model and Parameterization
where $\mu$ indexes all independent constraints in the network, including both bond-stretching terms $(i,j)\in\mathcal{L}$ and bond-bending terms $(i,j,k)\in\mathcal{A}$. The associated extensions are $e_{ij}=\hat{\mathbf{e}}_{ij}\cdot(\mathbf{u}_{i}-\mathbf{u}_{j})$ and $e_{ijk}=\theta_{ijk}-\theta^{0}_{ijk}$. Stacking all extensions into a vector $\mathbf{e}$ and displacements into $\mathbf{u}$ defines the rigidity matrix $R$ through $\mathbf{e}=R\mathbf{u}$. In our implementation, the same $R$ used for Maxwell–Calladine counting also enters the energetic dynamical matrix once local stiffness and prestress are specified, consistent with the generalized rigidity analysis of Rocks and Mehta [24]. The potential energy and dynamical matrix then follow as:
$V=\tfrac{1}{2}\,\mathbf{u}^{T}(R^{T}KR)\,\mathbf{u}\\$ (3)
$D=M^{-1/2}\,R^{T}K\,R\,M^{-1/2}$ (4)
Here $M$ is the diagonal mass matrix constructed from the atomic masses $m_{i}$, and $K$ are the bond and angle spring constants obtained from UFF4MOF parameters, with atom types assigned according to local coordination environments. Explicitly,
$K_{ij}$ $=664.12\,\frac{Z_{i}Z_{j}}{r_{ij}^{3}}$ (5)
$K_{ijk}$ $=\beta\;\frac{Z_{i}Z_{k}}{r_{ik}^{5}}\,\Big[\,3\,r_{ij}r_{jk}(1-\cos^{2}\theta_{0})-r_{ik}^{2}\cos\theta_{0}\Big],$ (6)
where $\beta=664.12/(r_{ij}r_{jk})$ and $Z_{\ell}$ are the UFF effective charges. We focus on these terms as they provide the dominant constraints governing framework rigidity. The UFF equilibrium distances are used as rest lengths; across all linear bonds they exceed the geometric bond lengths by 5–7% on average (see Appendix B).
The bonding network itself is determined using the CrystalNN[20] algorithm implemented in pymatgen[19]. A tunable cutoff parameter $\tau$ controls which coordination shells are included, as shown in Fig. 2. In the main dataset analysis we adopt the plateau $\tau$ identified by CrystalNN, which provides a chemically sensible minimal network, but in Fig. 5 we explore the effect of reducing $\tau$ for UiO-66, showing that modest additions of longer-range constraints systematically lift many of the baseline zero modes into soft, finite-frequency vibrations.
@@ II.2 Phonon Dispersions and Localization Analysis
which yields the $\mathbf{k}$-dependent dynamical matrix $D(\mathbf{k})$. Diagonalizing $D(\mathbf{k})$ gives the phonon spectrum $\omega^{2}(\mathbf{k})$, which we compute along the Setyawan–Curtarolo [26] high-symmetry paths in the Brillouin zone.
@@
Refer to caption(A)(B)(C)
Figure 4: Normalized distributions of the local Maxwell index $\nu_{i}$ for the three representative MOFs in Fig. 3. Positive $\nu_{i}$ indicates locally under-constrained environments, $\nu_{i}\approx 0$ near-isostatic environments, and negative $\nu_{i}$ locally over-constrained environments. ABIXOZ (A) exhibits a comparatively uniform, over-constrained distribution, whereas IKEBUV01 (B) and UiO-66 (C) show strong heterogeneity with under-constrained hydrogen-rich environments coexisting with a near-isostatic framework backbone (dominated by heavier atoms) UiO-66 additionally exhibits a strongly over-constrained metal-node population at large negative $\nu_{i}$.
@@ II.3 Dataset and Implementation Details
We apply this method to 5,682 structures from the CoRE MOF 2019 database [4]. Primitive cells are used whenever available to reduce redundancy. CIF [9] parsing, neighbor detection, constraint assembly, and phonon analysis were automated to enable large-scale screening. To validate our method, we benchmarked the rigidity code against crystalline silicon, shown in Appendix C, and against UiO-66, whose detailed analysis is presented in the main text. Full implementation details and all input files are provided in Appendix B and in the publicly available repository referenced therein.
@@
Figure 5: (A) Benchmark phonon dispersion calculated via GULP, reproduced from Ref. [11]. (B) Corresponding result from our rigidity-based model ($N_{0}=238$) using the default $\tau$ setting. (C) Decreasing the cutoff to $\tau=0.5$ introduces auxiliary constraints, reducing the zero mode count to $N_{0}=215$. (D) Further tuning to $\tau=0.035$ eliminates all zero modes ($N_{0}=0$). The progression from (B) to (D) demonstrates how assigning finite stiffness to long-range auxiliary bonds lifts zero modes into soft, collective bands, progressively filling the low-frequency spectrum relative to the benchmark in (A).
@@
Figure 6: Rigidity statistics across the CoRE MOF dataset. (A) Scatter plot of $\nu$ for 500 representative MOFs, with the blue dashed line marking the isostatic threshold ($\nu=3$). (B) Histogram of the Maxwell–Calladine index density ($\nu/N_{s}$) for 5,182 MOFs at $\Gamma$. The majority of frameworks are formally over-constrained yet cluster near the isostatic point. This concentration is dominated by accidental modes, which place many MOFs on the verge of mechanical instability.
@@ III Results
Our large-scale analysis of 5,682 structures from the CoRE MOF 2019 database shows that most frameworks are formally over-constrained at $\Gamma$. Figure 6 summarizes this dataset-level behavior in two complementary views. The histogram shows that the Maxwell–Calladine index density $\nu/N_{s}$ is typically negative, indicating an excess of constraints relative to degrees of freedom. However, rather than being broadly distributed, the population concentrates strongly near the isostatic reference $\nu/N_{s}=3/N$, which lies close to zero for large unit cells. The scatter plot shows that this near-isostatic concentration is widespread across the dataset rather than being driven by a small number of outliers.
Partitioning the statistics by zero-mode character shows that the dominant contribution to this near-isostatic peak comes from accidental modes. The largest share arises from accidental modes that are not confined to hydrogen yet also do not substantially involve metal-node motion. Hydrogen-only accidental modes form a secondary contribution, while accidental modes that significantly move metal atoms are comparatively rare.
This sharp concentration indicates a universal mechanical trend: the self-assembly of porous frameworks drives many MOFs toward the brink of mechanical instability. In our dataset, this marginal behavior is dominated by accidental geometric modes, which arise when the generic-geometry assumption behind Maxwell–Calladine counting is violated. The local geometry of metal clusters and linkers can create special alignments that render some constraints redundant, producing additional zero modes despite an overall excess of constraints. In contrast, frameworks classified as generic are shifted toward more negative $\nu/N_{s}$ values, consistent with conventional stability from independent constraints. To systematically characterize these behaviors, we classify frameworks into three regimes based on the relationship between $\nu$ and the vibrational spectrum:
@@
  • Geometrically Unstable ($\nu<3$, $N_{0}>3$): Symmetry creates constraint redundancy, producing accidental nontrivial mechanisms despite being formally over-constrained.
@@
Structures with $\nu>3$ are classified as under-constrained and are stabilized only by weak long-range couplings. In practical terms, generically rigid frameworks are robust to small distortions, whereas geometrically unstable and singular isostatic frameworks can be pushed into rigid behavior by modest geometric changes. We illustrate these regimes using the representative MOFs ABIXOZ, IKEBUV01, and UiO-66 in Figs. 3 and 4: Fig. 3 shows phonon dispersions together with inverse participation ratio (IPR) maps labeled by the dominant atomic species in each eigenvector, while Fig. 4 resolves the corresponding atomic-scale constraint balance via the local Maxwell index distributions $\nu_{i}$. This local index is defined for each atom i as:
$\nu_{i}=3-r_{i}$ (8)
where $r_{i}$ is the local constraint density. By partitioning the global constraints $\mu$ defined in Eq. (2), we assign:
$r_{i}=\frac{1}{2}n_{i,\mathcal{L}}+\frac{1}{3}n_{i,\mathcal{A}}$ (9)
where $n_{i,\mathcal{L}}$ is the number of bonds in $\mathcal{L}$ connected to atom $i$, and $n_{i,\mathcal{A}}$ is the number of bond angles in $\mathcal{A}$ in which the atom participates. This partitioning ensures that the local indices sum to the global index, $\sum\nu_{i}=\nu$.
ABIXOZ represents the generic rigidity regime. It is strongly over-constrained ($\nu/N_{s}=-1.89$) and contains no non-translation zero modes at $\Gamma$. Consistent with this, its dispersion in Fig. 3(3) is smooth and well ordered, and its IPR profile is predominantly delocalized with a stable dominant-species signature. The corresponding local-index distribution is comparatively uniform, consistent with a framework whose constraints are largely independent and whose rigidity is well predicted by global counting.
IKEBUV01 exemplifies the singular isostatic regime. At the mechanical critical point ($\nu=3$), its stability is symmetry-enforced rather than generic: it contains twelve nontrivial zero modes at $\Gamma$ that are exactly canceled by twelve states of self stress. In Fig. 3(3), this marginality appears as low-frequency asymmetries and nearly flat branches. The IPR map shows that the soft and zero-like motion at $\Gamma$ is overwhelmingly dominated by hydrogen, indicating that the accidental zero-mode subspace is largely confined to linker-associated, peripheral degrees of freedom rather than involving the load-bearing backbone. The local rigidity distribution in Fig. 4 mirrors this separation: hydrogen is concentrated at positive $\nu_{i}$, while the heavier framework atoms cluster near $\nu_{i}\approx 0$, consistent with a near-isostatic backbone coexisting with locally under-constrained hydrogen-rich environments. Despite being mechanically marginal by symmetry, this composition suggests that many of the zero modes are likely mechanically benign internal flexibilities.
UiO-66 lies in the geometrically unstable regime. Despite being globally over-constrained ($\nu/N_{s}=-1.14$), it exhibits 238 zero modes at $\Gamma$, indicating accidental geometric modes produced by constraint dependencies. In this zero-mode subspace, motion is dominated by hydrogen (about $83\%$ participation), with a substantial carbon component (about $16\%$) and only minor oxygen participation (about $1\%$), indicating that the soft degrees of freedom primarily reside on ligand-associated hydrogen and carbon environments rather than on the metal–oxygen backbone. In Fig. 3(3), these appear as bands that flatten near $\Gamma$, signaling compliant internal degrees of freedom within an otherwise over-constrained network; the associated IPR patterns show strong hydrogen participation at low frequency together with mixed-character modes at higher frequency. Figure 4 provides the structural origin: alongside a near-isostatic carbon–oxygen backbone, UiO-66 contains a strongly over-constrained metal-node population at very negative $\nu_{i}$, highlighting pronounced heterogeneity in local rigidity.
@@
To confirm the geometric origin of the UiO-66 zero modes, we systematically introduce auxiliary long-range bonds by tuning the cutoff parameter $\tau$ (Fig. 5). As $\tau$ decreases, the constraint density increases; assigning finite stiffness to these auxiliary bonds lifts the zero modes to finite frequencies, producing soft collective branches that progressively fill the low-frequency spectrum.
@@
Taken together, these case studies show that global constraint counting alone cannot predict rigidity in porous crystalline materials. ABIXOZ behaves as a generically rigid network, whereas IKEBUV01 and UiO-66 derive their mechanical character from accidental modes and constraint dependencies that are only apparent when global indices are complemented by spectral (IPR) and local ($\nu_{i}$) diagnostics.